OpenStudy (anonymous):

Which function has a vertex on the y-axis? a) f(x) = (x – 2)2 b)f(x) = x(x + 2) c)f(x) = (x – 2)(x + 2) d)f(x) = (x + 1)(x – 2)

OpenStudy (dtan5457):

That means which ever one's x coordinate for the vertex is 0, will have it's vertex on the y axis.

OpenStudy (dtan5457):

Try calculating the last 2 using (p+q)/2 for the x coordinate p,q=x intercepts

OpenStudy (anonymous):

ok one moment @dtan5457

OpenStudy (anonymous):

ok i got lost im not sure how to do it @dtan5457

OpenStudy (dtan5457):

p,q=x axis intercepts f(x) = (x – 2)(x + 2) p=2 q=-2 f(x) = (x + 1)(x – 2) p=-1 q=2

OpenStudy (dtan5457):

use (p+q)/2 and which ever one turns out to be 0 is your answer

OpenStudy (anonymous):

im not sure how to use the (p+q)/2 form @dtan5457

OpenStudy (dtan5457):

I just gave you the p and q values. just plug them in

OpenStudy (anonymous):

wait so it would be(2-2) (-2+2) @dtan5457

OpenStudy (anonymous):

@butterflydreamer

OpenStudy (anonymous):

@pitamar

OpenStudy (anonymous):

Remember the last problem? you have to find the vertex position of each one of those functions. Remember that if the function has 2 roots, then the vertex will be in their middle. if it has one root then this is the vertex. (you don't have here cases without roots). The root is basically a point where the function gives a value of 0 ( f(x) = 0 ). Can you tell for what 'x' values the first function gives a value of 0?

OpenStudy (anonymous):

so instead im adding 0 like this f(o)=(0-2)(0+2) @pitamar

OpenStudy (anonymous):

Not exactly. you basically wanna do something of the kind $$f(x)=0\implies (x-2)2 = 0 $$ and find what value x will make it true. Those values of 'x' that function value is 0 for are the 'roots' of the functions.

OpenStudy (anonymous):

so it would be 2? @pitamar

OpenStudy (anonymous):

Exactly. that is the single root of this function =) Not other value of 'x' will make this equation true. That means that the vertex is at x=2. You can find the y value for that vertex if you want, but the question is: `What function has a vertex on the y-axis?` The y-axis is basically at x=0. This vertex is not in x=0, so this function cannot be the answer

OpenStudy (anonymous):

so the answer is not c? @pitamar

OpenStudy (anonymous):

The answer is C =) The roots are 2 and -2 and the middle of those values is indeed 0.

OpenStudy (anonymous):

okay gotcha thanks :) @pitamar

OpenStudy (anonymous):

Np =)